A truck’s wheel has a diameter of 1m and cycles 26 times. How far does the truck move to the nearest meter?
a) 35m
b) 82m
c) 89cm
d) 109m
e) 111m

Answer by Theo(13342)

diameter of 1 meter means a radius of .5 meters.

circumference of the wheel is equal to 2*pi*r = 2*pi*.5 = pi = 3.141592654 meters.

1 cycle of the wheel rolling is equal to the circumference of the wheel.

26 cycles means the wheel has rolled the length of its circumference 26 times.

That equals to:

3.141592654 * 26 = 81.68140899 meters of rolling.

82 meters is the closest to 81.68140899 meters.

Your answer is selection b.